OpenStudy (anonymous):

Circlular metal plate for a machine has radius r. Larger plate with radius r+2 is to be usedin another part of the machine. Write an expression in simplest form for the sum of the areas of the two circles. my answers is: r^2+2r it that right?

OpenStudy (amistre64):

\[\pi r^2 + \pi(r+2)^2 \rightarrow \pi(r^2 +r^2 +4r +4)\]\[\rightarrow \pi(2r^2 +4r +4) = 2\pi(r^2 +2r +2)\]

OpenStudy (amistre64):

you had a part of it ;)

OpenStudy (anonymous):

so the anwser should be 2(r^2+2r+2)

OpenStudy (amistre64):

real close there.... you left out the "pi": 2 pi(r^2 +2r +2)